OpenStudy (anonymous):

I know the answer, but I need to know why its right.. Angela & Neil are going to the movies. They each bought a medium popcorn, and Neil got a small soft drink. Angela had a $5 gift certificate to put toward the cost, which came to $27.90. A movie ticket cost $10 each, and a medium popcorn costs $5.50. What was the cost for the soft drink?

OpenStudy (anonymous):

The answer is $1.90 , but why ?!

OpenStudy (anonymous):

So, 2 people going movie ticket is $10 each so 2 x 10 = 20 add on the popcorn 20 + 5.50x 2 = 20 + 11 = 31 (times 2 because there is 2 popcorns one for each) angela has a 5 gift certificate, so take away 5 so we get 31 - 5 = 26 Now the total cost was 27.90, so the difference is the price of the drink which is 27.90 - 26 = 1.90 which is reasonable for a small drink in a cinema